# How Do Companies Manage the Environmental Sustainability of AI? An Interview Study About Green AI Efforts and Regulations

### v1

@@ Abstract
Index Terms:
artificial intelligence, environmental sustainability, green AI, AI adoption, AI regulations, industry interviews
@@ I Introduction
The capabilities and impact of artificial intelligence (AI) [1] have increased drastically over the last decade, leading to both excitement and worry worldwide. A turning point in AI development occurred in November 2022 when OpenAI introduced a prototype of their chatbot ChatGPT [2], a Generative Pre-trained Transformer (GPT) based on a Large Language Model (LLM). Two months later, in January 2023, the Generative AI (GenAI) tool reached over 100 million monthly active users, making ChatGPT the fastest-growing consumer application in history [3]. However, training machine learning (ML) models like this and using them in complex applications at scale consumes vast amounts of energy [4], thereby contributing substantially to carbon emissions.
@@
The already achieved and further expected benefits of AI have led to an exponential increase in ML model size, bringing the environmental implications of AI to light [5]. For example, ChatGPT is built on GPT-4 foundation models containing over 1.76 trillion parameters, which is about 5,000 times as much as the Deep Neural Networks created in the early 2010s. As a result, the associated energy consumption, and thus the carbon impact have increased dramatically [6]. Chien et al. [7] estimated that ChatGPT-like chatbots produce 13 million kg of CO2 per year, with an inference of 11 million prompts per hour. Training these models requires over 50 million kWh, which can have a substantial negative impact on the environment [5]. As GenAI is becoming the fastest-growing technology in history, these emission rates are concerning.
@@
Reducing and remediating climate change induced by green house gas emissions is critical and requires immediate and decisive action. One of the key challenges in mitigating global warming is the transition towards more sustainable business, also known as the Green Transition [8]. On the level of the United Nations (UN), the European Union (EU), or related regulatory bodies, this is often connected to Environmental Social Governance (ESG), i.e., the sustainability aspects that organizations are assessed on [9]. Over the last decade, there has been an increasing emphasis on Corporate Social Responsibility (CSR) among organizations, leading to increased interest in ESG reporting [10]. For example, in January 2023, the Corporate Sustainability Reporting Directive (CSRD) entered into force [^1] https://eur-lex.europa.eu/eli/dir/2022/2464/oj, an EU regulation that establishes ESG reporting requirements for organizations. This new legislation aims to improve the accessibility of ESG reporting, increasing trust and transparency in the sustainability reports of companies and organizations [11]. Accordingly, sustainable development is expected to become a more important movement within companies globally.
@@
Similarly, the EU AI Act [^2] https://eur-lex.europa.eu/eli/reg/2024/1689/oj, the first European legislation on AI to ensure the responsible development and usage of AI, entered into force in August 2024. With the AI Act, the priority is to guarantee that AI-based software is safe, transparent, traceable, non-discriminatory, and environmentally friendly [12]. One of its key requirements is transparency and provision of information, i.e., companies have to become more transparent about how AI-based software is created and used, which may also help to assess its carbon emissions. The AI Act requires record keeping, i.e., automatic logging during system operation, to ensure traceability throughout the application of AI-based software [13]. It also mentions environmental sustainability directly, albeit less forcefully. For example, the AI Act refers to the principles of the EU’s High-Level Expert Group for AI (one of them is societal and environmental well-being) and article 95 requires the EU AI Office to facilitate codes of conduct for voluntary application that include assessing and minimizing the environmental impact of AI. All of this is meant to encourage companies to promote green and sustainable AI, aligning with the European Green Deal [14].
@@ II Background & Related Work
As the significant energy demand for AI became evident, Schwartz et al. [15] introduced the distinction between Red AI, which prioritizes the highest accuracy without considering energy efficiency, and Green AI, which aims to minimize its environmental impact while still achieving accurate results. Recently, Green AI has emerged as an active research area, as highlighted in a literature review by Verdecchia et al. [16] that identified 98 articles on the topic. More than 75% of these studies have been published since 2020. The review categorizes the Green AI literature into 13 main topics, with the most prominent being monitoring, hyperparameter tuning, model benchmarking, deployment, and model comparison. The authors suggest that interview studies could help understand how AI practitioners are currently addressing AI’s environmental impact. By synthesizing findings from our interviews with industry practitioners, we aim to provide a comprehensive understanding of the current landscape and to identify areas for future research in minimizing AI’s environmental footprint.
@@
Research on the negative environmental impact of AI. The energy consumption of machine learning models has been studied extensively in the last decade [17, 18, 19]. For example, already in 2016, Li et al. [20] conducted a comprehensive evaluation of the energy efficiency of training frameworks for deep Convolutional Neural Networks (CNNs). They presented a detailed study of the energy required for training using popular CNNs for image classification in computer vision, including a thorough analysis of different types of neural network layers. Beyond energy, Strubell et al. [4] analyzed the carbon impact of training their own state-of-the-art models. They concluded that developing and training such models incurs significant financial and environmental costs. Financially, the costs come from the hardware and electricity required, and environmentally, they result from the carbon footprint required to fuel modern tensor processing hardware.
@@
More recently, researchers examined the negative impact of Large Language Models (LLMs) on the environment, from the cost of training models to inference, highlighting the enormous recent and expected growth of AI. Wu et al. [5] explained how the carbon footprint of LLMs is largely due to scaled-up inference, despite the substantial energy demands for their training. As another example, Luccioni et al. [21] discussed the carbon footprint of the BLOOM model, a 176-billion-parameter language model. Using the Life Cycle Assessment (LCA) methodology, they evaluated the holistic environmental impact of the model. Taking a broader perspective, Luccioni et al. [22] provided a systematic comparison of the ongoing inference costs of AI models. They found that the most energy- and carbon-intensive tasks are those that generate new content, such as text generation, text summarization, or image generation. Some studies also estimated the carbon footprint of specific model architectures, such as GPT-3, Gopher, and OPT [23, 24, 25]. These studies focus primarily on the CO2 emissions produced by LLMs. However, the negative environmental impact of AI extends beyond CO2 emissions. For example, George et al. [26] investigated AI’s water usage and showed that GPT-3 consumed 700,000 liters of water during its training phase alone. The authors call for urgency to address this concern, as freshwater scarcity is a global problem [27].
@@
Tools for monitoring the negative environmental impact of AI. Another relevant research direction is the development of tools to measure or estimate the negative impacts of LLMs on the environment. For example, Budennyy et al. [28] introduced Eco2AI, an open-source package to track the energy consumption and CO2 emissions of AI models. Several similar open-source libraries are available to monitor CO2 emissions during the training of AI models [29, 30, 31]. Moreover, Li et al. [32] presented formulas to estimate the water footprint of AI models, encouraging organizations to consider water usage when developing, training, or using AI. Additionally, tools have been introduced to predict the carbon footprint of new neural networks before they have been trained. One early example is the Machine Learning Emissions Calculator by Lacoste et al. [33]. A more recent calculator is LLMCarbon by Faiz et al. [34], which is an end-to-end carbon footprint projection model designed for LLMs.
@@
Strategies for reducing the negative environmental impact of AI. Many studies have been conducted to investigate strategies to reduce the energy consumption and carbon footprint of AI models and systems [16]. The research community has begun to translate this knowledge into reusable design decisions, such as architectural tactics [35]. Another example is the study by Yarally et al. [36], which examined the effect of Bayesian optimization during hyperparameter tuning on the energy consumption of model training. They also investigated the complexity of AI models and concluded that the overall energy consumption of training can be halved by reducing the network complexity. Hyperparameter tuning and model complexity have increasingly become topics of interest in recent years [37, 38]. Additionally, research has focused on improving the environmental sustainability of model inference. For example, Li et al. [39] described how users can optimize their use of LLMs to minimize inference rates and reduce carbon emissions.
@@
AI adoption. Radhakrishnan and Chattopadhyay [40] conducted an in-depth analysis of 45 articles to identify the main theories and frameworks on AI adoption, including factors that facilitate, hinder, and determine the rate of diffusion in AI adoption in organizations. They showed that the most dominant theories used to study AI adoption are Technology, Organization, and Environment (TOE) and Diffusion of Innovation (DOI). At the organizational level, identified key barriers to AI adoption were digital maturity, trust, skill base, and privacy laws. However, the study does not cover AI sustainability, e.g., its energy consumption or carbon footprint.
@@
Governance, laws, and regulations. Recently, policymakers, practitioners, and academics have increasingly advocated for the development of global AI governance. In a paper called “Digital Sovereignty, Digital Expansionism, and the Prospects for Global AI Governance”, Roberts et al. [41] argue that, while global governance initiatives to manage AI technology risks seem promising, they doubt the efficacy of such initiatives in practice. They caution that aggressive pushes for digital sovereignty of individual countries may negatively impact such global governance initiatives, especially regarding the complex relationships between the EU, US, and China. Moreover, Birkmann et al. [11] studied the impact of the CSRD on large companies in Germany. Through 14 interviews with communication experts of these companies, they explored how the CSRD has influenced CSR communication. They concluded that it remains uncertain whether the CSRD will lead to standardized CSR communications across different companies.
@@ III Study Design
More specifically, we relied on a qualitative approach using semi-structured interviews [42], allowing for an in-depth exploration of this topic. Using semi-structured interviews provides rich and detailed results and also enables probing and discussing information that emerges from the participants’ responses, leading to more in-depth information to be collected [43]. We used the Interview Protocol Refinement (IPR) framework by Castillo-Montoya [44], which consists of four phases to systematically develop an interview protocol: 1) ensuring interview questions align with research questions, 2) constructing an inquiry-based conversation, 3) receiving feedback on interview protocols, and 4) piloting the interview protocol. To ensure transparency and replicability, we make our interview artifacts available on Zenodo. [^3] https://doi.org/10.5281/zenodo.14934325 This includes the interview preamble to recruit participants, the consent form, the interview guide itself, the slides used during the interview, and the key (anonymized) answers given by the participants.
@@ III-A Sampling of Participants
We imposed several criteria that participants had to meet to be included in the study. First and most importantly, participants had to be employees of a company that actively uses AI-based software. Additionally, participants should be able to provide information about the adoption and use of their AI-based software, and about the potential considerations of environmental sustainability in this context. These criteria, while important for our RQs, led to a relatively small pool of potential participants, as employees may work with AI but are not always knowledgeable about how the company decides on the adoption of AI or what role sustainability plays [45].
To recruit participants, we mostly made use of convenience sampling [46], i.e., we shared the call for participation within our personal networks via email and social media, as well as the network of Accenture. Furthermore, the snowball technique was used to find more suitable interviewees for this research [47], e.g., we asked participants to further distribute the call within their network. Although such referrals are an effective method to recruit participants, the backgrounds of the referrers and their referrals can be quite similar, resulting in a sample with fairly homogeneous work experiences. Initially, we aimed to interview around 15 people. However, over the course of the 5-month project, it became increasingly difficult to recruit more participants, mostly because of the requirement to have sufficient knowledge about the AI adoption and Green AI initiatives within the company or team. Several practitioners were interested in participating but did not fully meet this criterion. Consequently, our final sample consisted of 11 interviewees from 10 different companies, which still provides suitable diversity and breadth.
@@
C3 Consulting >>> 100,000 P3 Managing Director Data & AI within Strategy & Consulting 18 IT
@@
C4 Semiconductors 10,000 - 50,000 P4 Data and AI Analyst <<< 1 IT
@@
C9 Electronics & Manufacturing >>> 100,000 P10 Software Architecture and Team Automation Lead 7 IT
C10 Consulting 10,000 - 50,000 P11 IT Consultant <<< 1 IT
@@ III-B Study Execution & Data Collection
Before interviewing our 11 participants, one pilot interview was conducted to fine-tune the interview protocol, which is the fourth and final phase of the used IPR framework [44]. This pilot helped to improve the clarity of the phrasing and the order of the questions for better flow. It also led to the addition of optional prompts to encourage more detailed responses.
@@ III-C Data Analysis
Once the transcripts were final, we applied thematic analysis [48]. We first removed irrelevant information from the transcripts, such as words that had to be repeated due to mispronunciation or noise, or statements that did not answer any interview questions and were irrelevant for our RQs. We then annotated the pre-processed data to create a codebook that structured our analysis. A popular qualitative analysis tool, MAXQDA, was used to analyze the pre-processed data further [49]. We created labels that aligned with our RQs, e.g., the RQ1 analysis (consideration of sustainability during AI adoption) led to labels about drivers or identified concerns. The codebook made it easier to merge labels that contained similar information and to create categories. Overall, this was a lengthy iterative process, with labels being continuously refined, merged, and split. As the last step, we elicited hierarchically ordered categories and characterized their relationships to each other.

### v2

@@ I Introduction
The capabilities and impact of artificial intelligence (AI) [flasinski_introduction_2016] have increased drastically over the last decade, leading to both excitement and worry worldwide. A turning point in AI development occurred in November 2022 when OpenAI introduced a prototype of their chatbot ChatGPT [openai_chatgpt_2024], a Generative Pre-trained Transformer (GPT) based on a Large Language Model (LLM). Two months later, in January 2023, the Generative AI (GenAI) tool reached over 100 million monthly active users, making ChatGPT the fastest-growing consumer application in history [wu2023brief]. However, training machine learning (ML) models like this and using them in complex applications at scale consumes vast amounts of energy [strubell2020energy], thereby contributing substantially to carbon emissions.
@@
The already achieved and further expected benefits of AI have led to an exponential increase in ML model size, bringing the environmental implications of AI to light [wu2022sustainable]. For example, ChatGPT is built on GPT-4 foundation models containing over 1.76 trillion parameters, which is about 5,000 times as much as the Deep Neural Networks created in the early 2010s. As a result, the associated energy consumption, and thus the carbon impact have increased dramatically [kirkpatrick2023carbon]. chien2023reducing estimated that ChatGPT-like chatbots produce 13 million kg of CO2 per year, with an inference of 11 million prompts per hour. Training these models requires over 50 million kWh, which can have a substantial negative impact on the environment [wu2022sustainable]. As GenAI is becoming the fastest-growing technology in history, these emission rates are concerning.
@@
Reducing and remediating climate change induced by green house gas emissions is critical and requires immediate and decisive action. One of the key challenges in mitigating global warming is the transition towards more sustainable business, also known as the Green Transition [moodaley2023greenwashing]. On the level of the United Nations (UN), the European Union (EU), or related regulatory bodies, this is often connected to Environmental Social Governance (ESG), i.e., the sustainability aspects that organizations are assessed on [eccles2020social]. Over the last decade, there has been an increasing emphasis on Corporate Social Responsibility (CSR) among organizations, leading to increased interest in ESG reporting [aldogan2024determinants]. For example, in January 2023, the Corporate Sustainability Reporting Directive (CSRD) entered into force [^1] https://eur-lex.europa.eu/eli/dir/2022/2464/oj, an EU regulation that establishes ESG reporting requirements for organizations. This new legislation aims to improve the accessibility of ESG reporting, increasing trust and transparency in the sustainability reports of companies and organizations [birkmann2024csrd]. Accordingly, sustainable development is expected to become a more important movement within companies globally.
@@
Similarly, the EU AI Act [^2] https://eur-lex.europa.eu/eli/reg/2024/1689/oj, the first European legislation on AI to ensure the responsible development and usage of AI, entered into force in August 2024. With the AI Act, the priority is to guarantee that AI-based software is safe, transparent, traceable, non-discriminatory, and environmentally friendly [fidon2024dempster]. One of its key requirements is transparency and provision of information, i.e., companies have to become more transparent about how AI-based software is created and used, which may also help to assess its carbon emissions. The AI Act requires record keeping, i.e., automatic logging during system operation, to ensure traceability throughout the application of AI-based software [wagner2023navigating]. It also mentions environmental sustainability directly, albeit less forcefully. For example, the AI Act refers to the principles of the EU’s High-Level Expert Group for AI (one of them is societal and environmental well-being) and article 95 requires the EU AI Office to facilitate codes of conduct for voluntary application that include assessing and minimizing the environmental impact of AI. All of this is meant to encourage companies to promote green and sustainable AI, aligning with the European Green Deal [dyrhauge2024introduction].
@@ II Background & Related Work
As the significant energy demand for AI became evident, schwartz_green_2020 introduced the distinction between Red AI, which prioritizes the highest accuracy without considering energy efficiency, and Green AI, which aims to minimize its environmental impact while still achieving accurate results. Recently, Green AI has emerged as an active research area, as highlighted in a literature review by verdecchia2023systematic that identified 98 articles on the topic. More than 75% of these studies have been published since 2020. The review categorizes the Green AI literature into 13 main topics, with the most prominent being monitoring, hyperparameter tuning, model benchmarking, deployment, and model comparison. The authors suggest that interview studies could help understand how AI practitioners are currently addressing AI’s environmental impact. By synthesizing findings from our interviews with industry practitioners, we aim to provide a comprehensive understanding of the current landscape and to identify areas for future research in minimizing AI’s environmental footprint.
@@
Research on the negative environmental impact of AI. The energy consumption of machine learning models has been studied extensively in the last decade [sze2017hardware, yang2017designing, reagen2016minerva]. For example, already in 2016, li2016evaluating conducted a comprehensive evaluation of the energy efficiency of training frameworks for deep Convolutional Neural Networks (CNNs). They presented a detailed study of the energy required for training using popular CNNs for image classification in computer vision, including a thorough analysis of different types of neural network layers. Beyond energy, strubell2020energy analyzed the carbon impact of training their own state-of-the-art models. They concluded that developing and training such models incurs significant financial and environmental costs. Financially, the costs come from the hardware and electricity required, and environmentally, they result from the carbon footprint required to fuel modern tensor processing hardware.
@@
More recently, researchers examined the negative impact of Large Language Models (LLMs) on the environment, from the cost of training models to inference, highlighting the enormous recent and expected growth of AI. wu2022sustainable explained how the carbon footprint of LLMs is largely due to scaled-up inference, despite the substantial energy demands for their training. As another example, luccioni2023estimating discussed the carbon footprint of the BLOOM model, a 176-billion-parameter language model. Using the Life Cycle Assessment (LCA) methodology, they evaluated the holistic environmental impact of the model. Taking a broader perspective, luccioni2024power provided a systematic comparison of the ongoing inference costs of AI models. They found that the most energy- and carbon-intensive tasks are those that generate new content, such as text generation, text summarization, or image generation. Some studies also estimated the carbon footprint of specific model architectures, such as GPT-3, Gopher, and OPT [patterson2021carbon, berthelot2024estimating, cursaru2024controlled]. These studies focus primarily on the CO2 emissions produced by LLMs. However, the negative environmental impact of AI extends beyond CO2 emissions. For example, george2023environmental investigated AI’s water usage and showed that GPT-3 consumed 700,000 liters of water during its training phase alone. The authors call for urgency to address this concern, as freshwater scarcity is a global problem [gleick2021freshwater].
@@
Tools for monitoring the negative environmental impact of AI. Another relevant research direction is the development of tools to measure or estimate the negative impacts of LLMs on the environment. For example, budennyy2022eco2ai introduced Eco2AI, an open-source package to track the energy consumption and CO2 emissions of AI models. Several similar open-source libraries are available to monitor CO2 emissions during the training of AI models [henderson2020towards, dodge2022measuring, garcia2019estimation]. Moreover, li2025makingaithirstyuncovering presented formulas to estimate the water footprint of AI models, encouraging organizations to consider water usage when developing, training, or using AI. Additionally, tools have been introduced to predict the carbon footprint of new neural networks before they have been trained. One early example is the Machine Learning Emissions Calculator by lacoste2019quantifying. A more recent calculator is LLMCarbon by faiz2023llmcarbon, which is an end-to-end carbon footprint projection model designed for LLMs.
@@
Strategies for reducing the negative environmental impact of AI. Many studies have been conducted to investigate strategies to reduce the energy consumption and carbon footprint of AI models and systems [verdecchia2023systematic]. The research community has begun to translate this knowledge into reusable design decisions, such as architectural tactics [jarvenpaa_synthesis_2024]. Another example is the study by yarally2023uncovering, which examined the effect of Bayesian optimization during hyperparameter tuning on the energy consumption of model training. They also investigated the complexity of AI models and concluded that the overall energy consumption of training can be halved by reducing the network complexity. Hyperparameter tuning and model complexity have increasingly become topics of interest in recent years [panjapornpon2023explainable, kusumaraju24ecogen]. Additionally, research has focused on improving the environmental sustainability of model inference. For example, li2024toward described how users can optimize their use of LLMs to minimize inference rates and reduce carbon emissions.
@@
AI adoption. radhakrishnan2020determinants conducted an in-depth analysis of 45 articles to identify the main theories and frameworks on AI adoption, including factors that facilitate, hinder, and determine the rate of diffusion in AI adoption in organizations. They showed that the most dominant theories used to study AI adoption are Technology, Organization, and Environment (TOE) and Diffusion of Innovation (DOI). At the organizational level, identified key barriers to AI adoption were digital maturity, trust, skill base, and privacy laws. However, the study does not cover AI sustainability, e.g., its energy consumption or carbon footprint.
@@
Governance, laws, and regulations. Recently, policymakers, practitioners, and academics have increasingly advocated for the development of global AI governance. In a paper called “Digital Sovereignty, Digital Expansionism, and the Prospects for Global AI Governance”, roberts2023digital argue that, while global governance initiatives to manage AI technology risks seem promising, they doubt the efficacy of such initiatives in practice. They caution that aggressive pushes for digital sovereignty of individual countries may negatively impact such global governance initiatives, especially regarding the complex relationships between the EU, US, and China. Moreover, birkmann2024csrd studied the impact of the CSRD on large companies in Germany. Through 14 interviews with communication experts of these companies, they explored how the CSRD has influenced CSR communication. They concluded that it remains uncertain whether the CSRD will lead to standardized CSR communications across different companies.
@@ III Study Design
More specifically, we relied on a qualitative approach using semi-structured interviews [Hove2005], allowing for an in-depth exploration of this topic. Using semi-structured interviews provides rich and detailed results and also enables probing and discussing information that emerges from the participants’ responses, leading to more in-depth information to be collected [bhattacherjee2012social]. We used the Interview Protocol Refinement (IPR) framework by castillo2016preparing, which consists of four phases to systematically develop an interview protocol: 1) ensuring interview questions align with research questions, 2) constructing an inquiry-based conversation, 3) receiving feedback on interview protocols, and 4) piloting the interview protocol. To ensure transparency and replicability, we make our interview artifacts available on Zenodo. [^3] https://doi.org/10.5281/zenodo.14934325 This includes the interview preamble to recruit participants, the consent form, the interview guide itself, the slides used during the interview, and the key (anonymized) answers given by the participants.
@@ III-A Sampling of Participants
We imposed several criteria that participants had to meet to be included in the study. First and most importantly, participants had to be employees of a company that actively uses AI-based software. Additionally, participants should be able to provide information about the adoption and use of their AI-based software, and about the potential considerations of environmental sustainability in this context. These criteria, while important for our RQs, led to a relatively small pool of potential participants, as employees may work with AI but are not always knowledgeable about how the company decides on the adoption of AI or what role sustainability plays [smith2024clinicians].
To recruit participants, we mostly made use of convenience sampling [Baltes2022], i.e., we shared the call for participation within our personal networks via email and social media, as well as the network of Accenture. Furthermore, the snowball technique was used to find more suitable interviewees for this research [biernacki1981snowball], e.g., we asked participants to further distribute the call within their network. Although such referrals are an effective method to recruit participants, the backgrounds of the referrers and their referrals can be quite similar, resulting in a sample with fairly homogeneous work experiences. Initially, we aimed to interview around 15 people. However, over the course of the 5-month project, it became increasingly difficult to recruit more participants, mostly because of the requirement to have sufficient knowledge about the AI adoption and Green AI initiatives within the company or team. Several practitioners were interested in participating but did not fully meet this criterion. Consequently, our final sample consisted of 11 interviewees from 10 different companies, which still provides suitable diversity and breadth.
@@
C3 Consulting >> 100,000 P3 Managing Director Data & AI within Strategy & Consulting 18 IT
@@
C4 Semiconductors 10,000 - 50,000 P4 Data and AI Analyst << 1 IT
@@
C9 Electronics & Manufacturing >> 100,000 P10 Software Architecture and Team Automation Lead 7 IT
C10 Consulting 10,000 - 50,000 P11 IT Consultant << 1 IT
@@ III-B Study Execution & Data Collection
Before interviewing our 11 participants, one pilot interview was conducted to fine-tune the interview protocol, which is the fourth and final phase of the used IPR framework [castillo2016preparing]. This pilot helped to improve the clarity of the phrasing and the order of the questions for better flow. It also led to the addition of optional prompts to encourage more detailed responses.
@@ III-C Data Analysis
Once the transcripts were final, we applied thematic analysis [cruzes2011recommended]. We first removed irrelevant information from the transcripts, such as words that had to be repeated due to mispronunciation or noise, or statements that did not answer any interview questions and were irrelevant for our RQs. We then annotated the pre-processed data to create a codebook that structured our analysis. A popular qualitative analysis tool, MAXQDA, was used to analyze the pre-processed data further [maxqda]. We created labels that aligned with our RQs, e.g., the RQ1 analysis (consideration of sustainability during AI adoption) led to labels about drivers or identified concerns. The codebook made it easier to merge labels that contained similar information and to create categories. Overall, this was a lengthy iterative process, with labels being continuously refined, merged, and split. As the last step, we elicited hierarchically ordered categories and characterized their relationships to each other.